OpenStudy (anonymous):

What value of x makes this equation true? 2(x – 5) – 7 = 12 a. 19/2 b. -5 c. 7 d. -15

Directrix (directrix):

2(x – 5) – 7 = 12 Distribute the 2 over the (x-5) to get: 2x - 10 - 7 = 12 Question: What does -10 - 7 = @iNeedHomeWorkHelp

Directrix (directrix):

Take a guess. I am not allowed to give out answers without input from the Asker.

OpenStudy (anonymous):

-17. Then what? :) d thanks for your help!

Directrix (directrix):

That is correct. So picking up from: 2x - 10 - 7 = 12 we get 2x - 17 = 12 Adding 17 to both sides of the equation gives: 2x - 17 + 17 = 12 + 17 2x = 12 + 17 ----> What is 12 + 17 @iNeedHomeWorkHelp

Directrix (directrix):

@iNeedHomeWorkHelp --> What is 12 + 17 ?

OpenStudy (anonymous):

29.

Directrix (directrix):

Then, 2x = 12 + 17 and 2x = 29 and x = 29/2 --> Did you write the correct option for A? @iNeedHomeWorkHelp